### v1

@@ 2.1 Problem setup
Let $\bm{\zeta}_{t}=(\mu_{t},\sigma_{t}^{2})$ be parameters of the assumed DGP, $\bm{x}_{t-1}=\{r_{t-K},r_{t-K+1},\dots,r_{t-1}\}$ be a $K$-length input sequence [^2] For illustrative purposes, we have stated that the sequence only contains returns $r_{t}$. However, as discussed in Section 3.2, we also include squared returns $r_{t}^{2}$ as part of the input sequence. using returns up to $t-1$ and $y_{t-1}=r_{t}$ be forward one period return. The training dataset is comprised of $\mathcal{D}_{t}=\{(\bm{x}_{q-1},y_{q-1})|q\in\operatorname{\mathbb{N}}:q\leq t\}$ input-output pairs [^3] Note that at each portfolio selection period $t$, the training set can at most contain data up to $t-1$ as we have not yet observed $r_{t+1}$. and is essentially a set of sequences formed with a $K$-length sliding window and their corresponding regression targets. Our goal is to forecast $y_{t}$ (which corresponds to $r_{t+1}$). At each $t$, the investor’s goal is to solve the optimization problem [^4] For clarity, the case of a single asset is shown. At each $t$, there are $N$ assets and the dataset is typically in a $t\times N$ layout. It is easy to see the generalization of Equation 2 over $N$ assets, where the average loss is calculated over $(t-K-1)\times N$ instances.,
$\bm{\theta}_{t}=\operatorname*{\arg\min}_{\bm{\theta}^{*}}-\sum_{q=K}^{t-1}\log\mathrm{p}(y_{q}|F(\bm{x}_{q};\bm{\theta}^{*})),$ (2)
@@
where $F(\bm{x};\bm{\theta})$ is a neural network with input $\bm{x}$ and parameters $\bm{\theta}$, $\bm{\theta}=\bigcup_{\ell=1}^{L}\{\bm{W}^{(\ell)},\bm{b}^{(\ell)}\}$ is the set of network weights and biases and, in this context, $\mathrm{p}(y|F(\bm{x};\bm{\theta}))$ is the likelihood of observing $y$ based on the outputs of neural network $F(\cdot;\cdot)$ and the assumed marginal distribution. In other words, the investor is concerned with recovering the parameters $\hat{\bm{\zeta}}_{t}=(\hat{\mu}_{t},\hat{\sigma}_{t}^{2})\coloneqq F(\bm{x}_{t};\bm{\theta}_{t})$ that are most likely to have generated the observed data. In this setup, $\hat{\sigma}_{t}^{2}$ can be interpreted as an estimate of aleatoric uncertainty and is an estimate of the contemporaneous variance of the DGP at time $t$.
@@
There are two parts to this problem. The first part concerns UQ specifically for time-series that exhibit volatility clustering and is the primary focus of this work. The second part concerns advancing methods of UQ across general applications. In Appendix A.1, we show that our proposed approach can still benefit non-time-series problems in spite of it being designed to deal with a series of data points indexed in time order and exhibiting volatility clustering.
2.2 Related work
Recent advances in neural network UQ, such as Ensemble and Evidential, have focused on outputting parameters of the assumed data distribution. As these works were originally proposed for non-time-series problems, in discussing these works, we have left out time index $t$ but note that in our motivating application, variables are indexed by $t$ (e.g., the assumed DGP in Equation (1)). The neural networks are trained using procedures similar to maximum likelihood estimation. In Ensemble (Lakshminarayanan et al.,, 2017), regression target $y$ is assumed to be drawn from $y\sim\operatorname{N}(\mu,\sigma^{2})$, where $\mu$ is the forecast of $y$ and $\sigma^{2}$ models aleatoric uncertainty. The output layer of the neural network is modified to output $\bm{\zeta}=(\mu,\sigma^{2})$, and the network is trained using the Gaussian NLL. As this formulation is incapable of quantifying epistemic uncertainty, Lakshminarayanan et al., (2017) used an ensemble of neural networks with randomly initialized weights to provide an empirical estimate of epistemic uncertainty. Addressing this, Amini et al., (2020) proposed to place an evidential prior, the NIG distribution, on the model parameters $\mu,\sigma^{2}$ of the Normal data distribution:
@@ 2.2 Related work
NIG $\text{NIG prior}:\mu$ $\sim\operatorname{N}(\gamma,\sigma^{2}\nu^{-1}),\quad\sigma^{2}\sim\operatorname{InvGam}(\alpha,\beta),$ (3)
where $\mu$ is assumed to be drawn from a Normal prior distribution with unknown mean $\gamma$ and scaled variance $\sigma^{2}\nu^{-1}$, $\nu$ is a scaling factor for $\sigma^{2}$, and shape $\alpha>1$ and scale $\beta>0$ parameterize the IG distribution [^5] Time index $t$ has been omitted for brevity and legibility. Note that variables in this section are indexed by time for each asset: $\{y_{t},r_{t},\mu_{t},\sigma_{t}^{2},\gamma_{t},\nu_{t},\alpha_{t},\beta_{t}\}$.. We require $\alpha>1$ to ensure the mean of the marginal distribution is finite.
In this construct, parameters of the posterior distribution of $y$ is $\bm{\zeta}=(\gamma,\nu,\alpha,\beta)$. Epistemic uncertainty is reflected by the uncertainty in $\mu$, which is assumed be a fraction of $\sigma^{2}$ and is itself assumed to be drawn from an IG distribution. This fraction is controlled by $\nu$, which is learnt from the data and, in an abstract sense, varies according to the amount of information in the data. Parameter $\nu$ is interpreted as the number of virtual observations for the mean parameter $\mu$. In other words, $\nu$ virtual instances of $\mu$ are assumed to have been observed in determining the prior variance of $\mu$ (Jordan,, 2009; Amini et al.,, 2020).
For the NIG prior in 3, the marginal distribution of $\mu$ after integrating out $\sigma^{2}$ is a non-standardized Student’s t-distribution (denoted $\operatorname{St}$; Bernardo and Smith,, 2000),
@@
$=\operatorname{St}\left(\gamma,\frac{\beta}{\nu\alpha},2\alpha\right),$ (4)
using the fact that $\sigma^{2}\sim\operatorname{InvGam}(\alpha,\beta)$ corresponds to $\sigma^{-2}\sim\operatorname{Gam}(\alpha,\beta)$. Hence, assigning a $\operatorname{Gam}(\alpha,\beta)$ prior to precision $\sigma^{-2}$ in (Equation (3)) gives the Normal-Gamma (NG) prior and is equivalent to assigning $\operatorname{InvGam}(\alpha,\beta)$ to $\sigma^{2}$ which gives the NIG prior. The variance of this t-distribution is $\tfrac{\beta}{\nu(\alpha-1)}$. Predictions based on the NIG prior can be computed as (Amini et al.,, 2020),
$\text{Prediction}:\operatorname{E}[\mu]$ $=\gamma$
Aleatoric $\text{Aleatoric uncertainty}:\operatorname{E}[\sigma^{2}]$ $=\tfrac{\beta}{\alpha-1}$
The marginal variance $\operatorname{Var}[\mu]$ refers to the variance of the marginal t-distribution in (Equation (4)) for the NIG prior. We note that $\nu$ can also be interpreted as a factor that attributes uncertainty between aleatoric uncertainty ($\tfrac{\beta}{\alpha-1}$) and epistemic uncertainty ($\tfrac{\beta}{\nu(\alpha-1)}$). If $\nu=1$, then total uncertainty is evenly split between aleatoric and epistemic uncertainties.
Whilst not the focus of Amini et al., (2020), we note that epistemic uncertainty can be further decomposed approximately into uncertainties attributable to parameters $\mu$ and $\sigma^{2}$. Parameter $\mu|\sigma^{2}$ is normally distributed with $\operatorname{Var}[\mu|\sigma^{2}]=\sigma^{2}/\nu$ (from Equation (3)), and $\operatorname{E}[\sigma^{-2}]=\operatorname{E}[\tfrac{1}{\sigma^{-2}}]\approx\tfrac{1}{\operatorname{E}[\sigma^{-2}]}=\alpha/\beta$ (from the Gamma distribution of $\sigma^{-2}$). This leads to $\operatorname{Var}[\mu|\sigma^{2}]\approx\frac{\beta}{\nu\alpha}$,
@@
$\operatorname{Var}[y]=\frac{\beta}{\alpha-1}+\frac{\beta}{\nu(\alpha-1)}=\frac{\beta(1+\nu)}{\nu(\alpha-1)}.$ (8)
The corresponding NLL of Equation (7) is (Amini et al.,, 2020),
@@
Equation (9) mimics a Bayesian setup, granting classical neural networks the ability to estimate both epistemic and aleatoric uncertainty, and offers an intuitive interpretation of the model mechanics — due to uncertainty in the model parameters, the tails of the marginal likelihood are heavier than a Normal distribution. This has the effect of regularizing the network and provides an avenue of estimating epistemic uncertainty. As the distribution of asset returns has heavy tails (Cont,, 2001), we argue that the marginal t-distribution also provides a better fit of the data. The implementation is remarkably simple — Equation (9) replaces MSE as the loss function (for a regression problem) and the final layer of the network is replaced with a layer that simultaneously outputs four parameters of the marginal distribution. Clearly, modelling of $\gamma$ and $\nu$ by the neural network is direct as they correspond to mean and degrees of freedom of the t-distribution. By contrast, scale of the t-distribution in Equation (7) is modelled through a more complex structure ($\tfrac{\beta(1+\nu)}{\nu\alpha}$), which reflects the two sources of uncertainty in Equation (6) with two additional neural network outputs: $\alpha$ and $\beta$. They are the shape and scale parameters of the prior Gamma distribution for precision $\sigma^{-2}$ which describe distinct characteristics of epistemic uncertainty.
@@
As discussed in Section 1, aleatoric and epistemic uncertainties are difficult to disentangle. More recent works have questioned the accuracy of methods, such as Evidential, that directly estimate epistemic uncertainty through minimizing the NLL of an assumed marginal distribution. Bengs et al., (2023) argues that a strictly proper loss function for Evidential involves scoring against the distribution around each observation, which cannot possibly exist. Thus, there is no theoretical guarantee that the estimated epistemic uncertainty by Evidential is reliable. Moreover, Meinert et al., (2022) notes that Equation (7) is overparameterized, as it is possible to minimize Equation (9) irrespective of $\nu$, by: $\tfrac{\partial}{\partial\nu}\mathcal{L}_{\operatorname{NIG}}=0$, if $\beta\nu=\tfrac{1}{1+\nu^{-1}}$ and sending $\nu\to 0$. This is because Equation (7) is, by definition, a projection of the NIG distribution, and thus is unable to unfold all of its degrees of freedom unambiguously (Meinert et al.,, 2022). Through simulation data, Meinert et al., (2022) showed that over the course of neural network training, the estimated $\nu$ was related to speed of convergence. Thus, the estimated $\nu$, which controls the ratio of epistemic uncertainty to aleatoric uncertainty, may not be accurate. We note that this is also evident in Equation (7), as $\nu$ appears in both the numerator and denominator of the scale parameter of the t-distribution in the form of $1+\frac{1}{\nu}$. Thus, $\nu$ relates ambiguously to the the scale parameter of the t-distribution. Motivated by this observation, we propose a simpler formulation, which we detail in Section 3.1.
3 Uncertainty quantification under volatility clustering
3.1 Modelling forecast uncertainty using a scale mixture distribution

### v2

@@ 2.1 Problem setup
Let $\bm{\zeta}_{t}=(\mu_{t},\sigma_{t}^{2})$ be parameters of the assumed DGP, $\bm{x}_{t-1}=\{r_{t-K},r_{t-K+1},\dots,r_{t-1}\}$ be a $K$-length input sequence [^4] For illustrative purposes, we state that the sequence only contains returns $r_{t}$. However, as discussed in Section 2.4.3, we also include squared returns $r_{t}^{2}$ as part of the input sequence. using returns up to $t-1$ and $y_{t-1}=r_{t}$ be forward one period return. The training dataset is comprised of $\mathcal{D}_{t}=\{(\bm{x}_{q-1},y_{q-1})|q\in\operatorname{\mathbb{N}}:q\leq t\}$ input-output pairs [^5] Note that at each portfolio selection period $t$, the training set can at most contain data up to $t-1$ as we have not yet observed $r_{t+1}$. and is essentially a set of sequences formed with a $K$-length sliding window and their corresponding regression targets. Our goal is to forecast $y_{t}$ (which corresponds to $r_{t+1}$). At each $t$, the investor’s goal is to solve the optimization problem [^6] For clarity, the case of a single asset is shown. At each $t$, there are $N$ assets and the dataset is typically in a $t\times N$ layout. It is easy to see the generalization of Equation (2) over $N$ assets, where the average loss is calculated over $(t-K-1)\times N$ instances.,
$\bm{\theta}_{t}=\operatorname*{\arg\min}_{\bm{\theta}^{*}}-\sum_{q=K}^{t-1}\log\mathrm{p}\big{(}y_{q}|F(\bm{x}_{q};\bm{\theta}^{*})\big{)},$ (2)
@@
where $F(\bm{x};\bm{\theta})$ is a neural network with input $\bm{x}$ and parameters $\bm{\theta}$, $\bm{\theta}=\bigcup_{\ell=1}^{L}\{\bm{W}^{(\ell)},\bm{b}^{(\ell)}\}$ is the set of network weights and biases and $\mathrm{p}\big{(}y|F(\bm{x};\bm{\theta})\big{)}$ is the likelihood of observing $y$ based on the outputs of neural network $F(\cdot;\cdot)$ and the assumed marginal distribution. In other words, the investor is concerned with recovering the parameters $\hat{\bm{\zeta}}_{t}=(\hat{\mu}_{t},\hat{\sigma}_{t}^{2})\coloneqq F(\bm{x}_{t};\bm{\theta}_{t})$ that are most likely to have generated the observed data. In this setup, $\hat{\sigma}_{t}^{2}$ can be interpreted as an estimate of aleatoric uncertainty which is the contemporaneous variance of the DGP at time $t$.
@@
There are two parts to this problem. The first part concerns UQ specifically for time-series that exhibit volatility clustering and is the primary focus of this work. The second part concerns advancing methods of UQ across general applications. In Appendix B, we show that our proposed approach can still benefit non-time-series problems in spite of it being designed to deal with a series of data points indexed in time order and exhibiting volatility clustering.
2.2 Deep Ensemble method
Lakshminarayanan et al., (2017) proposed the Ensemble method which assumes that the regression target $y$ is drawn from a Normal distribution with mean $\mu$ and variance $\sigma^{2}$ denoted by $y\sim\operatorname{N}(\mu,\sigma^{2})$. The prediction of $y$ is taken at the mean level, that is, $\hat{y}=\mu$. They further introduced a Gaussian output layer in the network architecture, which simultaneously outputs $\bm{\zeta}=(\mu,\sigma^{2})$.
This model is implemented by minimizing the Gaussian NLL loss function,
$\mathcal{L}_{\operatorname{N}}(y|\mu,\sigma^{2})=\frac{1}{2}\log\left(2\pi\sigma^{2}\right)+\frac{(y-\mu)^{2}}{2\sigma^{2}},$ (3)
instead of the MSE in classical neural networks. Each network in the ensemble is trained using stochastic gradient descent. We remark that this model was originally proposed for non-time-series problems. Hence, we left out time index $t$ but in our motivating application, variables are indexed by $t$, similar to the DGP in Equation (1).
In this setup, $\sigma^{2}$ models aleatoric uncertainty but the Normal data distribution is incapable of quantifying epistemic uncertainty. Lakshminarayanan et al., (2017) addressed this by using an ensemble of neural networks with randomly initialized weights. Each network settles in a different local minima and produces different $\mu$ and $\sigma^{2}$ estimates for the same training dataset $\mathcal{D}_{t}$. The variance of $\hat{\mu}$ across the ensemble thus provides an empirically-driven proxy for epistemic uncertainty at a computational cost. Thus, Ensemble does not utilize statistical modelling to capture patterns in epistemic uncertainty.
2.3 Deep Evidential method
2.3.1 NIG prior distribution
Addressing the shortcomings of Ensemble, Amini et al., (2020) proposed to place an evidential prior, the NIG distribution, on the model parameters $\mu,\sigma^{2}$ of the Normal data distribution:
@@ 2.3.1 NIG prior distribution
NIG $\text{NIG prior}:\mu$ $\sim\operatorname{N}(\gamma,\sigma^{2}\nu^{-1}),\quad\sigma^{2}\sim\operatorname{IG}(\alpha,\beta),$ (4)
where $\mu$ is assumed to be drawn from a Normal prior distribution with unknown mean $\gamma$ and scaled variance $\sigma^{2}\nu^{-1}$, $\nu$ is a scaling factor for $\sigma^{2}$, and shape parameter $\alpha>1$ and scale parameter $\beta>0$ parameterize the Inverse Gamma (IG) distribution [^7] Time index $t$ has been omitted for brevity and legibility. Note that variables in this section are indexed by time for each asset: $\{y_{t},r_{t},\mu_{t},\sigma_{t}^{2},\gamma_{t},\nu_{t},\alpha_{t},\beta_{t}\}$.. As Evidential models epistemic uncertainty directly, no ensembling is necessary. Thus, compared to the Ensemble method, Evidential offers a computational cost advantage and a theoretically motivated epistemic uncertainty estimate. We require $\alpha>1$ to ensure the mean of the marginal distribution is finite.
In this construct, parameters of the posterior distribution of $y$ is $\bm{\zeta}=(\gamma,\nu,\alpha,\beta)$. Epistemic uncertainty is reflected by the uncertainty in $\mu$ and $\sigma^{2}$. The uncertainty in $\mu$ is assumed be a fraction $\nu^{-1}$ of $\sigma^{2}$ and $\sigma^{2}$ is assumed to be drawn from an IG distribution. The fraction $\nu^{-1}$ as controlled by $\nu$ is learnt from the data. In an abstract sense, $\nu$ varies according to the amount of information in the data and is interpreted as the number of virtual observations for the mean parameter $\mu$. In other words, $\nu$ number of virtual instances of $\mu$ are assumed to have been observed in determining the prior variance of $\mu$ (Jordan,, 2009; Amini et al.,, 2020).
Based on the NIG prior in Equation (4), predictions and sources of uncertainty are (Amini et al.,, 2020) [^8] Equation (5) is due to $\sigma^{2}\sim\operatorname{IG}(\alpha,\beta)$ with mean $\frac{\beta}{\alpha-1}$.,
$\text{Prediction}:\operatorname{E}[\mu]$ $=\gamma$
Aleatoric $\text{Aleatoric uncertainty}:\operatorname{E}[\sigma^{2}]$ $=\tfrac{\beta}{\alpha-1}$ (5)
Amini et al., (2020) derived these results using integration. We provide a more straight forward explanation using the marginal distribution of $\mu$ after integrating out $\sigma^{2}$ from the NIG prior. Bernardo and Smith, (2000) showed that this marginal distribution is a non-standardized Student’s t-distribution ($\operatorname{St}$),
@@
$=\operatorname{St}\left(\gamma,\frac{\beta}{\nu\alpha},2\alpha\right).$ (7)
Hence, the distribution of $\mu$ has mean $\gamma$, scale $\frac{\beta}{\nu\alpha}$ and degrees of freedom $2\alpha$ and so, $\operatorname{Var}[\mu]=\frac{\beta}{\nu\alpha}\times\frac{2\alpha}{2\alpha-2}=\frac{\beta}{\nu(\alpha-1)}$.
From Equations (5) and (6), $\nu$ can be interpreted as a factor that attributes uncertainty between aleatoric uncertainty ($\tfrac{\beta}{\alpha-1}$) and epistemic uncertainty ($\tfrac{\beta}{\nu(\alpha-1)}$). If $\nu=1$, then total uncertainty is evenly split between aleatoric and epistemic uncertainties. We also note that assigning a $\operatorname{Gamma}(\alpha,\beta)$ prior to precision $\sigma^{-2}$ in Bernardo and Smith, (2000) gives the Normal-Gamma (NG) prior and is equivalent to assigning $\operatorname{IG}(\alpha,\beta)$ to $\sigma^{2}$ which gives the NIG prior.
Apart from the direct explanation of Equation (6), we also enrich the interpretation of epistemic uncertainty from Amini et al., (2020) by decomposing it approximately into uncertainties attributable to parameters $\mu$ and $\sigma^{2}$. The Normal prior of $\mu|\sigma^{2}$ in Equation (4) has variance $\operatorname{Var}[\mu|\sigma^{2}]=\frac{\sigma^{2}}{\nu}\approx\frac{\beta}{\nu\alpha}$  [^9] Since $\sigma^{-2}\sim\operatorname{Gamma}(\alpha,\beta)$ with $\operatorname{E}[\sigma^{-2}]=\frac{\alpha}{\beta}$ such that $\operatorname{E}[\sigma^{2}]=\operatorname{E}[\tfrac{1}{\sigma^{-2}}]\approx\tfrac{1}{\operatorname{E}[\sigma^{-2}]}=\frac{\beta}{\alpha}$.. Hence, epistemic uncertainty can be decomposed into,
@@
$\operatorname{Var}[y]=\frac{\beta}{\alpha-1}+\frac{\beta}{\nu(\alpha-1)}=\frac{\beta(1+\nu)}{\nu(\alpha-1)},$ (10)
and is $(1+\nu)$ times the variance of $\mu$, the epistemic uncertainty. The corresponding NLL loss function using Equation (9) is (Amini et al.,, 2020),
@@
Equation (11) mimics a Bayesian setup, granting neural networks the ability to estimate both epistemic and aleatoric uncertainties, and offers an intuitive interpretation of the model mechanics — due to uncertainty in the model parameters, the tails of the marginal t-distribution are heavier than a Normal distribution. This has the effect of regularizing the network and provides an avenue of estimating epistemic uncertainty. To encourage low evidence (and hence high predictive uncertainty) in areas of $X$ with high forecast error, Amini et al., (2020) proposed to add a total evidence regularizer $\mathcal{L}_{\mathrm{R}}=|y-\gamma|(\nu+2\alpha)$ to $\mathcal{L}_{\operatorname{NIG}}$ (Equation (11)), where $\nu$ presents virtual observations for the mean and $2\alpha$ for the variance. As the distribution of asset returns has heavy tails (Cont,, 2001), we argue that the marginal t-distribution also provides better a fit for noisy financial time-series.
2.3.2 Shortcomings
Among the four parameters $(\gamma,\nu,\alpha,\beta)$, it is clear that the training of $\gamma$ in the neural network is direct as it corresponds to mean of the t-distribution in Equation (9). By contrast, the scale of t-distribution is modelled through a more complex construct ($\tfrac{\beta(1+\nu)}{\nu\alpha}$), which reflects the two sources of uncertainty in Equation (10) with two additional parameters: the shape ($\alpha$) and scale ($\beta$) parameters of the Gamma prior distribution for precision $\sigma^{-2}$. Bengs et al., (2023) argue that learning $\alpha,\beta$ in the hierarchical model (Equation (4)) requires $\mu,\sigma^{2}$ which are not directly observed. Thus, hierarchical models lack theoretical guarantee on the robustness of their estimated distributions.
@@ 2.3.2 Shortcomings
Moreover, Meinert et al., (2022) notes that Equation (9) is overparameterized, as it is possible to minimize Equation (11) irrespective of $\nu$ by sending $\nu\to 0$ (i.e., independent of the data), since $\tfrac{\partial}{\partial\nu}\mathcal{L}_{\operatorname{NIG}}=0$ along the path of $\beta\nu=\tfrac{1}{1+\nu^{-1}}$. This is because Equation (9) is, by definition, a projection of the NIG distribution, and thus is unable to unfold all of its degrees of freedom unambiguously (Meinert et al.,, 2022). Through simulation data, Meinert et al., (2022) showed that over the course of neural network training, the estimated $\nu$, which controls the ratio of epistemic uncertainty to aleatoric uncertainty, was related to speed of convergence. The results also show that the estimated $\nu$ may not be accurate. We note that this is also evident in Equation (9), as $\nu$ appears in both the numerator and denominator of ($\tfrac{\beta(1+\nu)}{\nu\alpha}$), the scale parameter of t-distribution in the form of $1+\frac{1}{\nu}$. Thus, $\nu$ relates ambiguously to the scale parameter of the t-distribution.
2.3.3 Architecture of neural network
Let $\bm{a}\in\operatorname{\mathbb{R}}^{H^{(I)}}$ be the input vector of the final layer with $H^{(I)}$ dimensions and $H^{(O)}$ be the dimension of the output layer. In the case of the NormalInverseGamma layer, $H^{(O)}=4$ as it outputs,
$\bm{\zeta}=\mathrm{O}(\bm{a};\bm{\theta})=\bm{a}^{\top}\cdot\bm{W}^{(O)}+\bm{b}^{(O)}$
[…]
As a further critique of the network architecture, we note that all four hyperparameters of Evidential are derived from the same latent representation outputted by the last hidden layer. Outputs of the NormalInverseGamma layer are linear transformations of a common input $\bm{a}$ (Equation (12)). The four hyperparameters can have vastly different scales (e.g., in our motivating application, $\gamma$ is in scale of 0.01, while $\nu$ is in scale of 10). We consider this feature to be a weakness of these approaches as the latent representation has to provide a sufficiently rich encoding to linearly derive all hyperparameters of the distribution. Nonetheless, successful applications of Evidential on real world datasets has led Meinert et al., (2022) to conclude that Evidential is a heuristic to Bayesian methods and may be appropriate for applications that aim to capture both aleatoric and epistemic uncertainties but do not demand an accurate distinction between them, such as our motivating application. Motivated by these observation, we propose a simpler formulation, which is more efficient in capturing both uncertainties. We detail the formulation in Section 2.4.1.
Figure 1: Our implementation of Evidential for cryptocurrency and U.S. equities forecasting. Output from the LSTM layers is fed into a single subnetwork, consisting of one or more blocks of fully connected and dropout layers. For illustrative purposes, we have shown the network with three LSTM layers and three hidden layer blocks. The actual number of hidden layers is subject to hyperparameter tuning (search range detailed in Appendix D). The NormalInverseGamma output layer (shaded in yellow) is a fully connected layer which outputs four hyperparameters of the marginal distribution. Softplus is applied to $\nu,\alpha,\beta$.
2.4 Proposed combined method
2.4.1 Scale mixture distribution